OpenStudy (anonymous):

Is a 79.59 a B OR C? In flvs

OpenStudy (♪chibiterasu):

0-59 = F 60 - 69 = D 70 - 74 = C- (usually passing) 75 - 79 = C+ (passing) 80 - 89 = B 90 - 100 = A

OpenStudy (anonymous):

So a C+

OpenStudy (meilendurcer):

there's a grade right next to that

OpenStudy (♪chibiterasu):

So yes, you passed your course. A C is a passing grade in FLVS. You need to get at least a D (a 60) on your finals to submit your final grade.

OpenStudy (♪chibiterasu):

Yes.

OpenStudy (anonymous):

Thnx, I was looking for a B-

OpenStudy (♪chibiterasu):

Yes it should say: http://prntscr.com/2c3hml If you want a B- you need to get at least 0.41 more points, which could be one simple re-submission :P

undeadknight26 (undeadknight26):

sometimes they round it and give you a B when you have a 79.6 or so.

OpenStudy (♪chibiterasu):

If you already submitted your final grade then that pretty much is your last grade for that segment of the course.

OpenStudy (anonymous):

I DID, noooo!!!

OpenStudy (♪chibiterasu):

In FLVS they don't round, my grade for U.S. History was 80.666666666 onward ._.

OpenStudy (♪chibiterasu):

They round to the nearest tenth I believe.

undeadknight26 (undeadknight26):

lol i got a 79.6 in OCVS and they rounded it...

OpenStudy (♪chibiterasu):

I mean hundreth.

OpenStudy (♪chibiterasu):

Yeah for me it's the nearest hundreth, so you got a 79.60

undeadknight26 (undeadknight26):

oh lol...when they issued my final grade i got an 80 and they counted that as a B.

OpenStudy (anonymous):

I wish I could give a medal to u @undeadknight26, look for a ? like 1 + 2, and I will give u it.

undeadknight26 (undeadknight26):

haha no im good don't need to go through all that trouble just to give me a medal :D

undeadknight26 (undeadknight26):

if medals were made of chocolate i would like one though :D

OpenStudy (anonymous):

ok cool thnx to everyone,

OpenStudy (♪chibiterasu):

I gave him a medal to be fair :3

undeadknight26 (undeadknight26):

tank you chib. i would @ you but i don't know that alt key code for the music sign off my head :(